form a polynomial with zeros:-2,2,3 and degree:3
Answer by ewatrrr(24785)


Hi
form a polynomial with zeros:-2,2,3 and degree:3
(x+2)(x-2)(x-3) = 0
(x+2)(x^2 - 5x + 6) = 0
x(x^2 - 5x + 6)+ 2(x^2 - 5x + 6)= 0
x^3 - 5x^2 + 6x + 2x^2 - 10x +12 = 0
x^3 - 3x^2 -4x + 12 = 0